OpenStudy (anonymous):

GEO1 HELP WILL GIVE MEDAL!

OpenStudy (anonymous):

Quadrilateral OPQR is inscribed inside a circle as shown below. What equation would be needed to solve for angle Q? What is the measure of angle Q? You must show all work and calculations to receive credit.

OpenStudy (anonymous):

@ganeshie8 @mathstudent55 @surjithayer @Ashleyisakitty @helpme1.2 @iheartfood @JoannaBlackwelder @Gabylovesyou @MrNood

OpenStudy (anonymous):

Hello pls help me out!

OpenStudy (gabylovesyou):

Try using the rule: O + Q = 180 R + P = 180 Explanation and proof: http://www.geom.uiuc.edu/~dwiggins/conj47.html O = (2x) Q = (2x + 4) O + Q = 180 This gives us the equation: 2x + (2x + 4) = 180 The measure of angle Q is the solution of that equation.

OpenStudy (anonymous):

in a cyclic quadrilateral opposite angles are supplementary. here <O+<Q=180

OpenStudy (anonymous):

WOW thank you!

OpenStudy (anonymous):

@surjithayer @Gabylovesyou so what exactly is Q tho?

OpenStudy (gabylovesyou):

have you tried to solve the equation ? :) look for x...

OpenStudy (anonymous):

find x and then put the value of x in <Q

OpenStudy (gabylovesyou):

yes ^ :)

OpenStudy (anonymous):

ya so X would be 4x right?

OpenStudy (gabylovesyou):

im not good at solving equations which is why i use my friend wolfram... http://www.wolframalpha.com/input/?i=2x+%2B+%282x+%2B+4%29+%3D+180+

OpenStudy (gabylovesyou):

x = ?

OpenStudy (anonymous):

How does that work?

OpenStudy (anonymous):

Oh 44

OpenStudy (gabylovesyou):

it says it there lol ;p you just input your equation and it solves it for you and YES correct :)

OpenStudy (anonymous):

Nice !

OpenStudy (anonymous):

4x+4=180 4x=180-4=176 x=176/4=44 <Q=?

OpenStudy (anonymous):

@surjithayer ya im doing it now

OpenStudy (anonymous):

x=176/4=44 would be 4

OpenStudy (anonymous):

So Q = 4?

OpenStudy (anonymous):

<Q=2 x+4=2*44+4=88+4=92

OpenStudy (anonymous):

lol sorry guys im dumb :p

OpenStudy (anonymous):

Can i tag both u guys on a next Question?

OpenStudy (gabylovesyou):

sure! :)